OpenStudy (rebecca1588):

Which equations in point-slope form are equations of the line that pass through the points (9, −2) and (−3, 6)?Choose exactly two answers that are correct. A. y+2=-2/3(x-9) B. y-9=-3/2(x+2) C. y+1=-3/2(x-4) D.y-6=-2/3(x+3)

pooja195 (pooja195):

First we need to find the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

hey buuuuddy! :D @rebecca1588 you need some help?

pooja195 (pooja195):

\[\LARGE \frac{6+2}{-3-9}=slope\]

OpenStudy (anonymous):

\[^^-He is-right-^^\]

pooja195 (pooja195):

Did you figure out the slope?

OpenStudy (rebecca1588):

y+2=-2/3(x-9) and y-6=-2/3(x+3) ? :)

pooja195 (pooja195):

Yep :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (rebecca1588):

Thanks 4 your help :D

OpenStudy (anonymous):

great job :D